OpenStudy (anonymous):

solve 3^(x + 4) = 50

OpenStudy (agreene):

log base 3 both sides and you get log _3 (x+4) = log _3 (50) from there you should be able to find your answer.

OpenStudy (anonymous):

i really dont know this, can you walk me through pls i will greatly aprreciate it. i need to understand this for a test, it isnt just a homework question

OpenStudy (anonymous):

3^(x+4) = 50 We have unknowns in the numerator. We use logarithms. Take a logarithm on both sides (could be base e or base 10 , it doesn't matter as long as you take the same base log on both sides. I will use base 10.

OpenStudy (anonymous):

log [ 3^(x+4)] = log(50) Now we use the rule log[x^n] = n log(x) , that is powers inside the logartithm can be brought out the front as factors. (x+4) log(3) = log(50) Now solve it for x (x+4) = log(50) / log(3)

OpenStudy (anonymous):

\[x= \frac{\log(50)}{\log(3)}-4 = -0.439\]

OpenStudy (anonymous):

If you sub into orginal equation it checks out.